A body of mass 100kg moves with a velocity of 20m/s in 2 minutes. Find the force of the body.
1
Expert's answer
2018-09-13T12:43:01-0400
m = 100kg
V = 20m/s
V0 = 0m/s
t = 2 min
F = ma
a = (V-V_0)/t= (20-0)/(2*60)=0.1667m/s^2
F = ma = 16.667 N